ihy:

Two particles A and B having different masses are projected from a tower with same speed. A is project vertically upward and B vertically downward. On reaching the ground... a) Velocity of A is greater than that of B b) Velocity of B is greater than that of A c) Both A and B attain the same velocity d) The particle with the larger mass attains higher velocity

breiadabrattzz:

Do you know the formula??

ihy:

@breiadabrattzz wrote:
Do you know the formula??
According to this there is none :sob:

breiadabrattzz:

It is

ihy:

@breiadabrattzz wrote:
It is
??

breiadabrattzz:

Let’s carefully analyze step by step

ihy:

@breiadabrattzz wrote:
Let’s carefully analyze step by step
Okay

breiadabrattzz:

Both A and B are projected from the same height (top of tower). Both are given the same initial speed (u). A goes vertically upward, while B goes downward. The acceleration on both is gravity (g) downward. Air resistance is neglected (since it’s not mentioned).

ihy:

@breiadabrattzz wrote:
Both A and B are projected from the same height (top of tower). Both are given the same initial speed (u). A goes vertically upward, while B goes downward. The acceleration on both is gravity (g) downward. Air resistance is neglected (since it’s not mentioned).
Alright

breiadabrattzz:

Particle A Initial velocity = –u (taking downward as positive direction). It goes up, stops momentarily, then falls down past the tower, gaining speed. Using the kinematic equation: v^2=u^2+2gh

ihy:

@breiadabrattzz wrote:
Particle A Initial velocity = –u (taking downward as positive direction). It goes up, stops momentarily, then falls down past the tower, gaining speed. Using the kinematic equation: v^2=u^2+2gh
Oh I got it!

ihy:

Is the answer C?

breiadabrattzz:

Yes

ihy:

YAY!

ihy:

Much appreciated :)

breiadabrattzz:

Your welcome.